OpenStudy (anonymous):

A triangular prism has a volume of 10 cm3. If each dimension of the prism is multiplied by four, what is the new volume of the prism? 320 cm3 80 cm3 160 cm3 640 cm3

OpenStudy (anonymous):

Volume of a triangular prism is given by 0.5 * base * height * length = 0.5 *b*h*l given 0.5*b*h*l = 10cm3 each dimension is now 4 times, 0.5 * 4b*4h*4l = (0.5*b*h*l) *4 *4 *4 Substitute 0.5*b*h*l = 10cm3 New Volume= 10*4*4*4= 640cm3

OpenStudy (anonymous):

Tiaph u r doing great to day

OpenStudy (anonymous):

Thanks!